### v1

# Some topological properties of the intrinsic volume metric

Abstract.
We prove that for fixed $j\in\{2,\ldots,d-1\}$, the $j$th intrinsic volume metric $\delta_{j}$ is not uniformly bounded, and the metric space $(\mathcal{K}_{j}^{d},\delta_{j})$ is incomplete. In particular, this proves that for $j\in\{2,\ldots,d-1\}$, the Hausdorff and $j$th intrinsic volume metrics do not induce the same topology on $\mathcal{K}_{j}^{d}$.
footnotetext: 2020 Mathematics Subject Classification: 52A27 (52A20, 52A39)footnotetext: Key words and phrases: convex body, intrinsic volume, metric, quermassintegral
1. Introduction and main results
Fix $d\in\mathbb{N}$. Let $\mathcal{K}^{d}$ denote the set of all convex, compact sets in $\mathbb{R}^{d}$. For $K\in\mathcal{K}^{d}$, we let $\dim(K)$ denote the dimension of the affine hull of $K$. For $j\in\{1,\ldots,d\}$, we set $\mathcal{K}_{j}^{d}:=\{K\in\mathcal{K}^{d}:\,\dim(K)\geq j\}$. For $K,L\in\mathcal{K}^{d}$, the Hausdorff distance $d_{H}(K,L)$ may be defined by
$d_{H}(K,L)=\inf\left\{\lambda\geq 0:\,K\subset L+\lambda B_{d},\,L\subset K+\lambda B_{d}\right\},$
where $B_{d}$ is the $d$-dimensional Euclidean unit ball centered at the origin. The symmetric difference metric is defined by $d_{S}(K,L)=\operatorname{vol}_{d}(K\triangle L)$, where $\operatorname{vol}_{d}$ denotes the $d$-dimensional volume functional and $K\triangle L=(K\setminus L)\cup(L\setminus K)$. It is well-known that the metric spaces $(\mathcal{K}^{d},d_{H})$ and $(\mathcal{K}_{d}^{d}\cup\{\varnothing\},d_{S})$ are complete. As explained in [BH-2024], in the latter case the empty set is appended to include limits of sequences of convex bodies $\{K_{i}\}\subset\mathcal{K}^{d}_{d}$ which converge in the Hausdorff metric to a set $K_{0}\in\mathcal{K}^{d}$ with $\dim(K_{0})<d$. More specifically, if $K_{i}\to K_{0}\in\mathcal{K}^{d}$ and $\dim(K_{0})<d$, then for such a sequence we have $d_{S}(K_{i},\varnothing)=\operatorname{vol}_{d}(K_{i})\to 0$, so $K_{i}\to\varnothing$ with respect to $d_{S}$. It is also known that on uniformly bounded subsets of $\mathcal{K}_{d}^{d}$, the metrics $d_{H}$ and $d_{S}$ are equivalent. For more background on the Hausdorff and symmetric difference metrics, see the article [ShephardWebster-1965] by Shephard and Webster.
For $j\in\{1,\ldots,d\}$ and $K,L\in\mathcal{K}_{j}^{d}$, the $j$th intrinsic volume metric $\delta_{j}(K,L)$ was defined in [BH-2024] by
(1) $\delta_{j}(K,L):=\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\int_{\operatorname{Gr}(d,j)}\operatorname{vol}_{j}\left((P_{H}K)\triangle(P_{H}L)\right)d\nu_{j}(H).$
Here and throughout the paper, $\operatorname{Gr}(d,j)$ is the Grassmannian manifold of all $j$-dimensional subspaces of $\mathbb{R}^{d}$, $\nu_{j}$ is the unique Haar probability measure on $\operatorname{Gr}(d,j)$, $P_{H}$ denotes the orthogonal projection of $\mathbb{R}^{d}$ into the subspace $H\in\operatorname{Gr}(d,j)$, and $\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]=\binom{d}{j}\frac{\operatorname{vol}_{d}(B_{d})}{\operatorname{vol}_{j}(B_{j})\operatorname{vol}_{d-j}(B_{d-j})}$ is the flag coefficient of Klain and Rota [Klain-Rota]. The intrinsic volume metric has found natural applications to the approximation of convex bodies by polytopes; see [BH-2024, BHK, HSW-2025].
It was conjectured in [BH-2024] that $d_{H}$ and $\delta_{j}$ induce the same topology on $\mathcal{K}_{j}^{d}$. In Open Question 2.2 of [BH-2024], they asked the following
Open Question 1.1 (uniform bound).
Let $j\in\{2,\dotsc,d-1\}$. If $\{K_{i}\}$ is a sequence in $\mathcal{K}_{j}^{d}$ that converges to $K\in\mathcal{K}_{j}^{d}$ with respect to $\delta_{j}$, does it follow that $\{K_{i}\}$ is uniformly bounded, that is, does there exist $R>0$ such that $\bigcup_{i\in\mathbb{N}}K_{i}\subset RB_{d}$?
As remarked in [BH-2024], a positive answer to this question would imply that a sequence $\{K_{i}\}$ which converges with respect to $\delta_{j}$ also converges with respect to $d_{H}$. Moreover, as noted in [BH-2024], for the special case $j=d$, this question was answered affirmatively by Shephard and Webster in [ShephardWebster-1965, Lemma 11], and for $j=1$ see [Vitale-1985, Theorem 3] and [Florian-1989, Theorem 2].
We show that, interestingly, when $2\leq j\leq d-1$, the answers to these questions are negative.
Theorem 1.2.
Let $2\leq j\leq d$ be given. There exists a sequence $\{K_{i}\}\subset\mathcal{K}_{j}^{d}$ such that
$\delta_{j}(K_{i},K)\to 0\quad\text{but}\quad\sup_{x\in K_{i}}\|x\|\to\infty\quad\text{as}\quad i\to\infty.$
Consequently, $d_{H}$ and $\delta_{j}$ induce different topologies on $\mathcal{K}_{j}^{d}$ if $2\leq j\leq d$.
Remark 1.3.
For $j=d$, note that $\delta_{d}=d_{S}$ is the symmetric difference metric. The nonequivalence in Theorem 1.2 is on $\mathcal{K}_{d}^{d}$ without a uniform boundedness restriction; on uniformly bounded families, the topologies induced by $d_{H}$ and $d_{S}$ do coincide.
Theorem 1.4.
Let $d\geq 3$ and $2\leq j\leq d-1$. The metric space $(\mathcal{K}_{j}^{d},\delta_{j})$ is incomplete.
It is thus natural to ask if appending the empty set to $\mathcal{K}_{j}^{d}$ will make $(\mathcal{K}_{j}^{d}\cup\{\varnothing\},\delta_{j})$ complete. While this is true for $j\in\{1,d\}$ (see [Florian-1989, Vitale-1985] for the case $j=1$), as we show below, it is still false for intermediate $j\in\{2,\ldots,d-1\}$.
Theorem 1.5.
Let $d\geq 3$ and $2\leq j\leq d-1$. The metric space $(\mathcal{K}_{j}^{d}\cup\{\varnothing\},\delta_{j})$ is incomplete.
Throughout the paper, we use $c,C,C_{1},C_{2}$, etc. to denote positive absolute constants whose value may change from line to line. If a positive constant depends on some parameter, the dependence will always be made explicit, e.g., $c(d)$ is a constant that depends only on $d$.
2. Proofs
2.1. Proof of Theorem 1.2
Let $E:=\operatorname{span}\{e_{1},\ldots,e_{j}\}\subset\mathbb{R}^{d}$ and fix a $j$-dimensional convex body $K\subset E$ (so $K\in\mathcal{K}_{j}^{d}$). Choose $x_{0}\in\operatorname{relint}(K)$ (relative to $E$) and a unit vector $u\in E$. Shrinking $\varepsilon$ if necessary, we may assume that $x_{0}+\varepsilon B_{u^{\perp}\cap E}^{j-1}\subset K$. For parameters $L>0$ and $\varepsilon>0$, define the following thin cylinder inside $E$:
(2) $N(L,\varepsilon):=\left\{x_{0}+tu+v:\,t\in[0,L],\,v\in\varepsilon B_{u^{\perp}\cap E}^{j-1}\right\}\subset E.$
Here $B_{u^{\perp}\cap E}^{j-1}:=\{x\in u^{\perp}\cap E:\,\|x\|\leq 1\}$. Now set
(3) $K(L,\varepsilon):=\operatorname{conv}\left(K\cup N(L,\varepsilon)\right)\subset E.$
Note that $\dim K(L,\varepsilon)=j$ and $K(L,\varepsilon)\in\mathcal{K}_{j}^{d}$ for all $L,\varepsilon>0$.
Fix $H\in\operatorname{Gr}(d,j)$. Let $u_{H}:=(P_{H}u)/\|P_{H}u\|$ when $P_{H}u\neq 0$, and decompose $H$ as the direct sum $H=\mathbb{R}u_{H}\oplus E_{H}$ where $E_{H}:=(P_{H}u)^{\perp}\cap H$. Let $\pi_{H}$ denote the orthogonal projection along $u_{H}$. Then for every $y\in E_{H}$, the fiber $(P_{H}(K(L,\varepsilon)))_{y}\setminus(P_{H}K)_{y}$ is either empty or is a single segment of length at most $L\|P_{H}u\|$, and this can only occur when $y\in\pi_{H}(P_{H}(\varepsilon B_{u^{\perp}\cap E}^{j-1}))$. Hence, with $\ell_{H}:=\|P_{H}u\|$ we get
$\operatorname{vol}_{j}\left([P_{H}(K(L,\varepsilon))]\triangle(P_{H}K)\right)\leq L\ell_{H}\operatorname{vol}_{j-1}\left(\pi_{H}(P_{H}(\varepsilon B_{u^{\perp}\cap E}^{j-1}))\right).$
Therefore, integrating over $H$ we obtain
$\delta_{j}(K(L,\varepsilon),K)$ $\leq\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\int_{\operatorname{Gr}(d,j)}L\ell_{H}\operatorname{vol}_{j-1}(P_{H}(\varepsilon B_{u^{\perp}\cap E}^{j-1}))\,d\nu_{j}(H).$
The orthogonal projection is 1-Lipschitz on $H$ and thus does not increase the (j1)(j-1)-dimensional volume. Hence,
$\operatorname{vol}_{j-1}(P_{H}(\varepsilon B_{u^{\perp}\cap E}^{j-1}))\leq\operatorname{vol}_{j-1}(\varepsilon B_{u^{\perp}\cap E}^{j-1})=\operatorname{vol}_{j-1}(B_{u^{\perp}\cap E}^{j-1})\varepsilon^{j-1}=\operatorname{vol}_{j-1}(B_{j-1})\varepsilon^{j-1}.$
Thus, using also $\ell_{H}\leq 1$, we get
$\delta_{j}(K(L,\varepsilon),K)$ $\leq L\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\operatorname{vol}_{j-1}(B_{j-1})\varepsilon^{j-1}=C(d,j)L\varepsilon^{j-1}$
where $C(d,j):=\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\operatorname{vol}_{j-1}(B_{j-1})$.
Now choose a sequence $\{L_{i}\}$ with $L_{i}\to\infty$, and set $\varepsilon_{i}:=L_{i}^{-2}$. For any fixed $j\geq 2$, we have
$\delta_{j}(K(L_{i},\varepsilon_{i}),K)\leq C(d,j)L_{i}\varepsilon_{i}^{j-1}=C(d,j)L_{i}^{1-2(j-1)}=C(d,j)L_{i}^{3-2j}\to 0\quad\text{as}\quad i\to\infty.$
and $\sup_{x\in K(L_{i},\varepsilon_{i})}\|x\|\to\infty$ as $i\to\infty$. Thus, we have proved the first assertion of the theorem with $K_{i}:=K(L_{i},L_{i}^{-2})\in\mathcal{K}_{j}^{d}$.
To prove the second assertion, consider the point $x_{i}:=x_{0}+L_{i}u\in N(L_{i},\varepsilon_{i})\subset K(L_{i},\varepsilon_{i})$. Let
$R_{K}:=\sup\{\|x\|:\,x\in K\}<\infty.$
Observe that by the reverse triangle inequality,
$\|x_{i}\|=\|x_{0}+L_{i}u\|\geq\|L_{i}u\|-\|x_{0}\|=L_{i}-\|x_{0}\|,$
and, for any $x\in K$,
$\|x_{i}-x\|\geq\|x_{i}\|-\|x\|\geq L_{i}-\|x_{0}\|-R_{K}.$
Therefore,
$d_{H}(K(L_{i},\varepsilon_{i}),K)\geq\operatorname{dist}(x_{i},K)\geq L_{i}-\|x_{0}\|-R_{K}\longrightarrow\infty\quad\text{as}\quad i\to\infty.$
$2.$ Proof of Theorem 1.4
Fix a $j$-dimensional subspace $E\in\mathrm{Gr}(d,j)$ and a unit vector $u\in E$. Given $x\in E$, $L>0$ and $\varepsilon>0$, define the following thin cylinder (a $j$-dimensional convex body inside $E$):
$N(x;L,\varepsilon):=\operatorname{conv}\left(x+[-L,L]u,x+\varepsilon B^{j-1}_{u^{\perp}\cap E}\right)\subset E,$
where $B^{j-1}_{u^{\perp}\cap E}$ is the Euclidean unit ball in $u^{\perp}\cap E$. For any $K\in\mathcal{K}^{d}_{j}$, set
$K^{+}(x;L,\varepsilon):=\operatorname{conv}(K\cup N(x;L,\varepsilon)).$
There exists a constant $C(d,j)>0$ (depending only on $d,j$) such that for all $K$, all $x\in E$, and all $L,\varepsilon>0$,
(4) $\delta_{j}(K^{+}(x;L,\varepsilon),K)\leq C(d,j)L\varepsilon^{j-1}.$
Indeed, for each $H\in\mathrm{Gr}(d,j)$, the $j$-dimensional volume in $H$ contributed by the added cylinder is bounded above by
$(\text{length of }P_{H}\text{-image of the segment})$ $\times((j-1)\text{-volume of the transverse cross-section})$
$\leq 2L\operatorname{vol}_{j-1}(B_{j-1})\varepsilon^{j-1},$
and we integrate this bound over $H$ (the projection is 11-Lipschitz on $E$). Multiplying both sides of this inequality by the fixed factor $\big[\!\begin{smallmatrix}d\\ j\end{smallmatrix}\!\big]$ and setting $C(d,j):=2\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\operatorname{vol}_{j-1}(B_{j-1})$, we obtain (4).
Next, we construct a $\delta_{j}$-Cauchy sequence whose projections escape to infinity. Fix any $K_{0}\in\mathcal{K}^{d}_{j}$ and let $E$ and $u$ be as above. Choose a sequence $\{L_{m}\}_{m\geq 0}$ with $L_{m}\to\infty$. Define $\varepsilon_{m}>0$ by
(5) $\varepsilon_{m}^{j-1}L_{m}=\frac{2^{-(m+1)}}{C(d,j)}.$
We inductively define
$K_{m+1}:=\operatorname{conv}(K_{m}\cup N(x_{m};L_{m},\varepsilon_{m})),$
where $x_{m}:=x_{0}+T_{m}u$ and $T_{m}\to\infty$ is arbitrary (e.g., $T_{m}=m$). Then for a.e. $H$ with $\|P_{H}u\|>0$, we have $\|P_{H}x_{m}\|>\|P_{H}u\|T_{m}-\|P_{H}x_{0}\|\to\infty$ as $m\to\infty$. Hence, for any prescribed $R_{m}\to\infty$, we eventually have $P_{H}x_{m}\in H\setminus B_{H}(0,R_{m})$. In particular, the projected cylinder $P_{H}N(x_{m};L_{m},\varepsilon_{m})$ lies outside $B_{H}(0,R_{m})$ for all sufficiently large $m$. By (4) and (5), we have
(6) $\delta_{j}(K_{m+1},K_{m})\leq C(d,j)L_{m}\varepsilon_{m}^{j-1}\ =2^{-(m+1)}.$
Hence $\sum_{m}\delta_{j}(K_{m+1},K_{m})<\infty$, and $\{K_{m}\}$ is $\delta_{j}$-Cauchy.
Suppose by way of contradiction that $\{K_{m}\}$ converges in the metric $\delta_{j}$ to some $K_{\infty}\in\mathcal{K}^{d}_{j}$:
$\delta_{j}(K_{m},K_{\infty})=\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\int_{\operatorname{Gr}(d,j)}\operatorname{vol}_{j}((P_{H}K_{m})\triangle(P_{H}K_{\infty}))\,d\nu_{j}(H)\longrightarrow 0\quad\text{as}\quad m\to\infty.$
Then with $f_{m}(H):=\operatorname{vol}_{j}((P_{H}K_{m})\triangle(P_{H}K_{\infty}))$, this implies $f_{m}\to 0$ in $L^{1}(H)$. By the definition of $\delta_{j}$, since $f_{m}\to 0$ in $L^{1}(H)$ there exists a subsequence (still denoted $\{f_{m}\}$) such that for $\nu_{j}$-almost every $H\in\mathrm{Gr}(d,j)$,
(7) $f_{m}(H)=\operatorname{vol}_{j}((P_{H}K_{m})\triangle(P_{H}K_{\infty}))\longrightarrow 0\quad\text{as}\quad m\to\infty.$
Next, we show that for a.e. $H$, each new cylinder contributes a fixed positive $j$-dimensional volume outside $B_{H}(0,R_{m})=\{y\in H:\,\|y\|\leq R_{m}\}$. We will need the following lemma, proved below.
Lemma 2.1.
Fix $E\in\mathrm{Gr}(d,j)$ and a unit vector $u\in E$. For $\nu_{j}$-almost every $H\in\mathrm{Gr}(d,j)$, the map $P_{H}|_{E}:E\to H$ has full rank $j$, and the linear map
$T_{H}:=\bigl(\pi_{H}\circ P_{H}\bigr)\big|_{u^{\perp}\cap E}:\ u^{\perp}\cap E\longrightarrow E_{H},$
where $E_{H}:=\{x\in H:\langle x,u_{H}\rangle=0\}$ and $u_{H}:=\frac{P_{H}u}{\|P_{H}u\|}$, is an isomorphism with (j1)(j-1)-Jacobian $J_{j-1}(T_{H})>0$.
Fix an $H$ that satisfies Lemma 2.1, and set
$\ell_{H}:=\|P_{H}u\|>0,\qquad b(H):=J_{j-1}(T_{H})\operatorname{vol}_{j-1}(B_{j-1})>0,\qquad c(H):=2\ell_{H}b(H)>0.$
Since $P_{H}$ is linear and $T_{H}$ is an isomorphism, the $H$-projection of the cylinder contains a rectangular block,
$P_{H}N(x_{m};L_{m},\varepsilon_{m})\ \supset\ \Big(P_{H}x_{m}+[-L_{m},L_{m}]\cdot P_{H}u\Big)\ \times\ \pi_{H}\!\Big(P_{H}\big(\varepsilon_{m}B^{\,j-1}_{u^{\perp}\cap E}\big)\Big),$
whose $j$-dimensional volume is at least
(8) $\operatorname{vol}_{j}\big(P_{H}N(x_{m};L_{m},\varepsilon_{m})\big)\geq 2L_{m}\|P_{H}u\|\varepsilon_{m}^{j-1}J_{j-1}(T_{H})\operatorname{vol}_{j-1}(B_{j-1})=c(H)\varepsilon_{m}^{j-1}L_{m}.$
By (5), the right-hand side equals $c(H)C(d,j)^{-1}2^{-(m+1)}$.
Since $P_{H}K_{m}\subset B_{H}(0,\rho_{m})$ for some $\rho_{m}<\infty$, choose $R_{m}>\rho_{m}$. Thus, for all large enough $m$, $P_{H}N(x_{m};L_{m},\varepsilon_{m})\subset H\setminus B_{H}(0,R_{m})$ and hence is disjoint from $P_{H}K_{m}$. Thus, since $P_{H}K_{m+1}=\operatorname{conv}(P_{H}K_{m}\cup P_{H}N(x_{m};L_{m},\varepsilon_{m}))$, and since we choose $R_{m}$ so that $P_{H}K_{m}\subset B_{H}(0,\rho_{m})\subset B_{H}(0,R_{m})$, for $\nu_{j}$-a.e. $H\in\operatorname{Gr}(d,j)$, there exists $m_{0}\in\mathbb{N}$ such that for all $m\geq m_{0}$,
$\big(P_{H}K_{m+1}\setminus P_{H}K_{m}\big)\cap\big(H\setminus B_{H}(0,R_{m})\big)\supset P_{H}N(x_{m};L_{m},\varepsilon_{m}).$
and
(9) $\begin{split}\operatorname{vol}_{j}\Big(\big(P_{H}K_{m+1}\setminus P_{H}K_{m}\big)\cap\big(H\setminus B_{H}(0,R_{m})\big)\Big)&\geq\operatorname{vol}_{j}\big(P_{H}N(x_{m};L_{m},\varepsilon_{m})\big)\\ &\geq c(H)C(d,j)^{-1}2^{-(m+1)}.\end{split}$
In the final step, we will obtain a contradiction using $L^{1}(H)$-convergence to a bounded projection. Fix the above $H$. The sets $P_{H}K_{\infty}$ are bounded in $H$, so there exists $R>0$ such that $P_{H}K_{\infty}\subset B_{H}(0,R)$. From (7), for the subsequence (which we still denote by $\{K_{m}\}$), we have
$\operatorname{vol}_{j}\big((P_{H}K_{m})\triangle(P_{H}K_{\infty})\big)\longrightarrow 0\quad\text{as}\quad m\to\infty.$
In particular, for this fixed $R$, $\operatorname{vol}_{j}((P_{H}K_{m})\setminus B_{H}(0,R))\to 0$ as $m\to\infty$. But $R_{m}\to\infty$, so eventually $R_{m}>R$, and then (9) yields a uniform positive lower bound
$\operatorname{vol}_{j}((P_{H}K_{m+1})\setminus B_{H}(0,R))\geq c(H)C(d,j)^{-1}2^{-(m+1)}>0$
for infinitely many $m$. This contradicts the fact that $\operatorname{vol}_{j}\big((P_{H}K_{m})\setminus B_{H}(0,R)\big)\to 0$ as $m\to\infty$. Therefore, no $K_{\infty}\in\mathcal{K}^{d}_{j}$ can be the $\delta_{j}$-limit of $\{K_{m}\}$, so $(\mathcal{K}^{d}_{j},\delta_{j})$ is incomplete.
Proof of Lemma 2.1.
We first show the statement on the full rank of $P_{H}|_{E}$. For $v\in E\setminus\{0\}$, the set $\mathcal{A}_{v}:=\{H\in\mathrm{Gr}(d,j):\,H\subset v^{\perp}\}$ identifies with $\mathrm{Gr}(d-1,j)$, a proper smooth submanifold of strictly lower dimension. Hence, $\nu_{j}(\mathcal{A}_{v})=0$. Let $\mathbb{S}(E):=\{v\in E:\|v\|=1\}$ denote the unit sphere in $E$ centered at the origin. Since $\mathbb{S}(E)$ is compact, it contains a countable dense subset $Q\subset\mathbb{S}(E)$. For $q\in Q$, define
$\mathcal{A}_{q}:=\left\{H\in\mathrm{Gr}(d,j):\,H\subset q^{\perp}\right\}.$
Each $\mathcal{A}_{q}$ is naturally identified with $\mathrm{Gr}(d-1,j)$, so $\dim\mathcal{A}_{q}=j(d-j-1)<j(d-j)=\dim\mathrm{Gr}(d,j)$, and therefore $\nu_{j}(\mathcal{A}_{q})=0$. If $H\in\mathrm{Gr}(d,j)$ is such that $H\subset v^{\perp}$ for some nonzero $v\in E$, then $v/\|v\|\in\mathbb{S}(E)\cap H^{\perp}$, a nonempty closed subset of $\mathbb{S}(E)$. Since $Q$ is dense in $\mathbb{S}(E)$, there exists $q\in Q\cap H^{\perp}$. Thus $H\subset q^{\perp}$, so $H\in\mathcal{A}_{q}$. This yields
$\{H\in\operatorname{Gr}(d,j):\,\exists v\in E\setminus\{0\}\text{ with }H\subset v^{\perp}\}\subset\bigcup_{q\in Q}\mathcal{A}_{q},$
which is a countable union of $\nu_{j}$-null sets and thus is a $\nu_{j}$-null set. Thus, for $\nu_{j}$-a.e. $H$, $E\cap H^{\perp}=\{0\}$ and $P_{H}|_{E}$ is bijective.
Next, we show that $T_{H}$ is (j1)(j-1)-isomorphism on $u^{\perp}\cap E$. Fix any $H$ in which $P_{H}|_{E}$ has full rank, so $P_{H}u\neq 0$ and $u_{H}$ is defined. By the first part, this set has $\nu_{j}$ measure 1. If $T_{H}w=0$ for some $w\in u^{\perp}\cap E$, then $P_{H}w\in\mathrm{span}\{u_{H}\}$, say $P_{H}w=\lambda u_{H}=\lambda P_{H}u/\|P_{H}u\|$ for some $\lambda\in\mathbb{R}$. Then $P_{H}(w-\beta u)=0$ where $\beta=\lambda/\|P_{H}u\|$, and since $w-\beta u\in E$ and $P_{H}|_{E}$ is injective, we get $w-\beta u=0$. Hence $0=\langle u,w\rangle=\beta\langle u,u\rangle=\beta$, so $w=0$. Therefore, $T_{H}$ is injective. Since $\dim(u^{\perp}\cap E)=\dim(E_{H})=j-1$, $T_{H}$ is an isomorphism, and its (j1)(j-1)-Jacobian is positive. ∎
2.3. Proof of Theorem 1.5
Fix a $j$-dimensional subspace $E\in\operatorname{Gr}(d,j)$ with $2\leq j\leq d-1$, a unit vector $u\in E$, and a nonempty compact convex set $K_{0}\subset E$ with $\dim K_{0}=j$. For $L>0$, $\varepsilon>0$ and a base point $x_{0}\in K_{0}$, define the thin cylinder
$N(x_{0};L,\varepsilon):=\operatorname{conv}\left(x_{0}+[-L,L]u,x_{0}+\varepsilon B^{j-1}_{u^{\perp}\cap E}\right),$
and set $K^{+}(x_{0};L,\varepsilon):=\operatorname{conv}(K_{0}\cup N(x_{0};L,\varepsilon))$. As in the proof of Theorem 1.4, for all $L,\varepsilon>0$,
(10) $\delta_{j}(K^{+}(x_{0};L,\varepsilon),K_{0})\leq C(d,j)L\varepsilon^{j-1}.$
where $C(d,j):=2\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\operatorname{vol}_{j-1}(B_{j-1})>0$.
We will construct a $\delta_{j}$-Cauchy sequence with uniformly positive distance to $\varnothing$. Let $a_{0}:=\delta_{j}(K_{0},\varnothing)$, which is positive since $K_{0}\neq\varnothing$. Choose any sequence $L_{m}\to\infty$, and define
(11) $\varepsilon_{m}:=\left(\frac{2^{-(m+1)}a_{0}}{4C(d,j)L_{m}}\right)^{\frac{1}{j-1}}.$
Define inductively $K_{m+1}:=\operatorname{conv}(K_{m}\cup N(x_{m};L_{m},\varepsilon_{m}))$, where the choice of the base point $x_{m}\in E$ will be specified later. Then by (10) and (11),
(12) $\delta_{j}(K_{m+1},K_{m})\leq C(d,j)L_{m}\varepsilon_{m}^{j-1}=\frac{a_{0}}{4}\cdot 2^{-(m+1)}.$
Hence, $\sum_{m}\delta_{j}(K_{m+1},K_{m})\leq a_{0}/4$, so $\{K_{m}\}$ is $\delta_{j}$-Cauchy. Moreover, by the triangle inequality,
$\left|\delta_{j}(K_{m+1},\varnothing)-\delta_{j}(K_{m},\varnothing)\right|\leq\delta_{j}(K_{m+1},K_{m})\leq\frac{a_{0}}{4}\cdot 2^{-(m+1)},$
so $\{\delta_{j}(K_{m},\varnothing)\}$ is Cauchy in $\mathbb{R}$ and thus converges to some $\alpha\in[0,\infty)$. We now show that $\alpha>0$. Note that for every $m$,
$\delta_{j}(K_{m},\varnothing)\geq\delta_{j}(K_{0},\varnothing)-\delta_{j}(K_{0},K_{m})\geq a_{0}-\sum_{i=0}^{m-1}\delta_{j}(K_{i+1},K_{i})\geq a_{0}-\frac{a_{0}}{4}=\frac{3}{4}a_{0}>0.$
Passing to the limit, we thus obtain
(13) $\alpha=\lim_{m\to\infty}\delta_{j}(K_{m},\varnothing)\geq\frac{3}{4}a_{0}>0.$
Therefore, $\{K_{m}\}$ does not converge to $\varnothing$ in $(\mathcal{K}_{j}^{d}\cup\{\varnothing\},\delta_{j})$.
Let $\mathscr{G}\subset\operatorname{Gr}(d,j)$ denote the set of “good” subspaces in Lemma 2.1. Then by Lemma 2.1, $\nu_{j}(\mathscr{G})=1$. For fixed $H\in\mathscr{G}$, define
$\ell_{H}:=\|P_{H}u\|>0,\quad b(H):=J_{j-1}(T_{H})\operatorname{vol}_{j-1}(B_{j-1})>0,\quad c(H):=2\ell_{H}b(H)>0.$
We now choose the base points so that projected mass appears arbitrarily far from the origin. Let $\{R_{m}\}_{m\geq 0}$ be any sequence with $R_{m}\to\infty$. Since $P_{H}|_{E}$ is onto, for each $m$ we can choose the base point $x_{m}\in E$ so that $P_{H}x_{m}$ lies sufficiently far in the $u_{H}-$direction to guarantee
(14) $P_{H}N(x_{m};L_{m},\varepsilon_{m})\subset H\setminus B_{H}(0,R_{m}).$
Since the projection is linear and $T_{H}$ is an isomorphism, the projected cylinder contains a rectangular block in $H$ whose $j$-volume is bounded below by
(15) $\operatorname{vol}_{j}\big(P_{H}N(x_{m};L_{m},\varepsilon_{m})\big)\geq 2L_{m}\|P_{H}u\|\varepsilon_{m}^{j-1}J_{j-1}(T_{H})\operatorname{vol}_{j-1}(B_{j-1})=c(H)\varepsilon_{m}^{j-1}L_{m}.$
By (11), the right-hand side equals $\frac{c(H)a_{0}}{4C(d,j)}\cdot 2^{-(m+1)}$. Moreover, since $P_{H}K_{m}$ is contained in some ball $B_{H}(0,\rho_{m})$ with $\rho_{m}<\infty$ (as by continuity, projections of compact sets are compact), choosing $R_{m}>\rho_{m}$ makes $P_{H}N(x_{m};L_{m},\varepsilon_{m})$ disjoint from $P_{H}K_{m}$, and from (14) and (15) we get the lower bound
(16) $\operatorname{vol}_{j}\Big(\big(P_{H}K_{m+1}\setminus P_{H}K_{m}\big)\cap\big(H\setminus B_{H}(0,R_{m})\big)\Big)\ \geq\frac{c(H)a_{0}}{4C(d,j)}\cdot 2^{-(m+1)}>0.$
Finally, we show that the sequence $\{K_{m}\}$ has no limit in $\mathcal{K}_{j}^{d}$. Suppose by way of contradiction that $K_{m}\to K_{\infty}$ in $\delta_{j}$ for some $K_{\infty}\in\mathcal{K}_{j}^{d}$. Following along the same lines as before, there exists a subsequence (again, not relabeled) such that for $\nu_{j}$-a.e. $H$,
(17) $\operatorname{vol}_{j}\big((P_{H}K_{m})\triangle(P_{H}K_{\infty})\big)\longrightarrow 0\quad\text{as}\quad m\to\infty.$
Fix the good $H\in\mathscr{G}$ chosen above, so that (17) holds for these $H$. Since $K_{\infty}$ is compact, $P_{H}K_{\infty}\subset B_{H}(0,R)$ for some $R>0$. Since $R_{m}\to\infty$, for all sufficiently large $m$ we have $R_{m}>R$, and hence
$\operatorname{vol}_{j}\big((P_{H}K_{m})\setminus B_{H}(0,R)\big)\longrightarrow 0.$
But (16) yields, for infinitely many $m$,
$\operatorname{vol}_{j}\big((P_{H}K_{m+1})\setminus B_{H}(0,R)\big)\geq\frac{c(H)a_{0}}{4C(d,j)}\cdot 2^{-(m+1)}>0,$
a contradiction. Therefore, $\{K_{m}\}$ has no limit in $\mathcal{K}_{j}^{d}$.
To conclude the proof, note that by (13), the sequence does not converge to $\varnothing$, and by the last step, it does not converge to any $K_{\infty}\in\mathcal{K}_{j}^{d}$ either. Hence, $(\mathcal{K}_{j}^{d}\cup\{\varnothing\},\delta_{j})$ is not complete. ∎
Department of Mathematics & Computer Science, Longwood University,
U.S.A.
E-mail address: hoehnersd@longwood.edu