### v2

# Some topological properties of the intrinsic volume metric

Abstract.
The purpose of this note is to derive certain basic, but previously unrecorded, topological properties of the intrinsic volume metrics $\delta_{1},\ldots,\delta_{d}$ on the space of convex bodies in $\mathbb{R}^{d}$. Our main results show that for every $2\leq j\leq d-1$, the topology induced by $\delta_{j}$ does not control the Hausdorff metric on the class of $j$-dimensional convex bodies; in particular, the condition $\delta_{j}(K_{n},K)\to 0$ does not imply uniform boundedness in the ambient space. Furthermore, for every $2\leq j\leq d-1$, the metric space $(K_{j}^{d},\delta_{j})$ is incomplete, and remains incomplete even after adjoining the empty set.
Our main results demonstrate that the intrinsic volume metric behaves in a fundamentally different way from the familiar Hausdorff and symmetric difference metrics. We describe the geometric mechanism that produces these phenomena and discuss implications for geometric tomography, metric stability theory and integral geometry.
footnotetext: 2020 Mathematics Subject Classification: 52A27 (52A20, 52A39)footnotetext: Key words and phrases: convex body, intrinsic volume, metric, quermassintegral
1. Introduction and main results
Metrics on the space of convex bodies are indispensable tools throughout convex geometry, geometric analysis and geometric tomography. The Hausdorff metric governs compactness phenomena and the Blaschke selection principle; the Banach–Mazur distance underlies asymptotic classification questions; and the symmetric difference metric is fundamental in shape optimization, stochastic geometry and random approximation theory.
The present paper concerns another natural family of metrics: those induced by the intrinsic volumes. For an integer $d\geq 2$ and a fixed $j\in\{1,\ldots,d\}$, the Kubota formula representation of the $j$th intrinsic volume of a convex body $K\subset\mathbb{R}^{d}$ is
$V_{j}(K)=\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\int_{\mathrm{Gr}(d,j)}\mathrm{vol}_{j}(P_{H}K)\,d\nu_{j}(H).$
Here and throughout the paper, $\operatorname{Gr}(d,j)$ is the Grassmannian manifold of all $j$-dimensional subspaces of $\mathbb{R}^{d}$, $\nu_{j}$ is the unique Haar probability measure on $\operatorname{Gr}(d,j)$, $P_{H}$ denotes the orthogonal projection of $\mathbb{R}^{d}$ into the subspace $H\in\operatorname{Gr}(d,j)$, and
$\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]:=\binom{d}{j}\frac{\operatorname{vol}_{d}(B_{d})}{\operatorname{vol}_{j}(B_{j})\operatorname{vol}_{d-j}(B_{d-j})}$
is the flag coefficient of Klain and Rota [Klain-Rota], where $B_{m}$ is the Euclidean unit ball in $\mathbb{R}^{m}$.
In view of Kubota’s formula, for convex bodies $K,L\subset\mathbb{R}^{d}$ and $j\in\{1,\ldots,d\}$, Besau and Hoehner [BH-2024] recently defined the $j$th intrinsic volume metric by
(1) $\delta_{j}(K,L)=\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\int_{\mathrm{Gr}(d,j)}\operatorname{vol}_{j}\left((P_{H}K)\triangle(P_{H}L)\right)d\nu_{j}(H).$
This metric measures the distance between two convex bodies via the average $j$-dimensional discrepancy of their orthogonal projections. Note that if $j=d$, then (1) reduces to the symmetric difference metric.
Projection-based distances of the form (1) arise naturally in the study of projection inequalities, Urysohn-type inequalities, stability problems, and reconstruction in geometric tomography. In applied contexts, $\delta_{j}$ quantifies differences between shapes when only lower-dimensional data are available, a situation which arises frequently in computerized tomography, stereology, and the analysis of high-dimensional data through random projections. In [BH-2024, BHK, HSW-2025], this metric was used to study the asymptotic best approximation of convex bodies by polytopes.
1.1. Motivation
A longstanding theme in metric geometry is to understand how different metrics interact: when they are equivalent, when they control one another, and when they yield fundamentally different topologies. In general, comparing the various metrics on convex bodies is a fundamental question that has been studied extensively; see, for example, [BH-2024, Florian-1989, Jin-Leng-Guo-2014-orlicz-metric, Groemer2000, ShephardWebster-1965, Vitale-1985].
With the recent introduction of the intrinsic volume metric in [BH-2024], it is thus natural to further investigate how this new metric compares to the other well-studied metrics, such as the Hausdorff and symmetric difference metric. Despite the central role of intrinsic volumes and their projection representations in convex geometry, many of the basic topological properties of the metrics $\delta_{j}$ have not yet been studied. The aim of this paper is to close these gaps while identifying several structural features that distinguish $\delta_{j}$ sharply from the Hausdorff and symmetric difference metrics.
Furthermore, understanding the topology induced by the intrinsic volume metric is important for several reasons, especially as it relates to applications:
  • Stability of geometric inequalities. Many inequalities involving intrinsic volumes, including Urysohn-type and Alexandrov–Fenchel inequalities, require sharp stability estimates. Determining whether $\delta_{j}$-convergence controls geometric degeneracy is helpful for formulating meaningful stability results.
  • Geometric tomography. Reconstruction algorithms often rely on comparing projection data. If $\delta_{j}$-convergence does not prevent “unbounded drift” in the Hausdorff metric, this limits identifiability and affects the choice of regularization.
  • Random approximations and stochastic geometry. Models of random polytopes or random projections approximate a target body in low-dimensional marginals. Whether such approximations converge to genuine convex bodies depends on the completeness and compactness properties of $(K_{j}^{d},\delta_{j})$.
1.2. Main results
Let $\mathcal{K}_{j}^{d}$ denote the set of $j$-dimensional convex bodies contained in some $j$-dimensional affine subspace of $\mathbb{R}^{d}$. Our first theorem shows that the intrinsic-volume metrics do not enforce any form of ambient boundedness.
Theorem 1.1 (Failure of uniform boundedness).
Let $2\leq j\leq d$ be given. There exists a sequence $\{K_{i}\}\subset\mathcal{K}_{j}^{d}$ such that
$\delta_{j}(K_{i},K)\to 0\quad\text{but}\quad\sup_{x\in K_{i}}\|x\|\to\infty\quad\text{as}\quad i\to\infty.$
Consequently, $d_{H}$ and $\delta_{j}$ induce different topologies on $\mathcal{K}_{j}^{d}$ if $2\leq j\leq d$.
This illustrates that $\delta_{j}$ captures only the average projection geometry of a body, not its global embedding in $\mathbb{R}^{d}$. This distinction is crucial in problems where ambient-space control is essential.
Our next results further establish that $\delta_{j}$ behaves quite differently from the Hausdorff and symmetric difference metrics from a metric-geometric point of view.
Theorem 1.2 (Incompleteness).
Let $d\geq 3$ and $2\leq j\leq d-1$. The metric space $(\mathcal{K}_{j}^{d},\delta_{j})$ is incomplete.
It is thus natural to ask if appending the empty set to $\mathcal{K}_{j}^{d}$ will make $(\mathcal{K}_{j}^{d}\cup\{\varnothing\},\delta_{j})$ complete. While this is true for $j\in\{1,d\}$ (see [Florian-1989, Vitale-1985] for the special case $j=1$), as we show below, it is still false for intermediate $j\in\{2,\ldots,d-1\}$.
Theorem 1.3 (Incompleteness with $\varnothing$ adjoined).
Let $d\geq 3$ and $2\leq j\leq d-1$. The metric space $(\mathcal{K}_{j}^{d}\cup\{\varnothing\},\delta_{j})$ is incomplete.
These negative results have both analytic and structural implications. They rule out direct use of fixed-point theorems, gradient-flow constructions and variational arguments based on completeness, and they clarify the limitations of projection-based metrics in quantifying geometric proximity.
Remark 1.4.
For $j=d$, note that $\delta_{d}=d_{S}$ is the symmetric difference metric. The nonequivalence in Theorem 1.1 is on $\mathcal{K}_{d}^{d}$ without a uniform boundedness restriction; on uniformly bounded families, the topologies induced by $d_{H}$ and $d_{S}$ do coincide.
The proofs of the main theorems are given in Section 3. The geometric idea in the proofs is to construct thin cylinders added far away to produce vanishing projection deviation but large Hausdorff drift, while the projection of cylinders inside “good” subspaces produces persistent mass arbitrarily far away, contradicting possible convergence.
2. Background and notation
Fix $d\in\mathbb{N}$. Let $\mathcal{K}^{d}$ denote the set of all convex, compact sets in $\mathbb{R}^{d}$. For $K\in\mathcal{K}^{d}$, we let $\dim(K)$ denote the dimension of the affine hull of $K$. For $j\in\{1,\ldots,d\}$, we set $\mathcal{K}_{j}^{d}:=\{K\in\mathcal{K}^{d}:\,\dim(K)\geq j\}$. For $K,L\in\mathcal{K}^{d}$, the Hausdorff distance $d_{H}(K,L)$ may be defined by
$d_{H}(K,L)=\inf\left\{\lambda\geq 0:\,K\subset L+\lambda B_{d},\,L\subset K+\lambda B_{d}\right\},$
where $B_{d}$ is the $d$-dimensional Euclidean unit ball centered at the origin. The symmetric difference metric is defined by $d_{S}(K,L)=\operatorname{vol}_{d}(K\triangle L)$, where $\operatorname{vol}_{d}$ denotes the $d$-dimensional volume functional and $K\triangle L=(K\setminus L)\cup(L\setminus K)$. It is well-known that the metric spaces $(\mathcal{K}^{d},d_{H})$ and $(\mathcal{K}_{d}^{d}\cup\{\varnothing\},d_{S})$ are complete. As explained in [BH-2024], in the latter case the empty set is appended to include limits of sequences of convex bodies $\{K_{i}\}\subset\mathcal{K}^{d}_{d}$ which converge in the Hausdorff metric to a set $K_{0}\in\mathcal{K}^{d}$ with $\dim(K_{0})<d$. More specifically, if $K_{i}\to K_{0}\in\mathcal{K}^{d}$ and $\dim(K_{0})<d$, then for such a sequence we have $d_{S}(K_{i},\varnothing)=\operatorname{vol}_{d}(K_{i})\to 0$, so $K_{i}\to\varnothing$ with respect to $d_{S}$. It is also known that on uniformly bounded subsets of $\mathcal{K}_{d}^{d}$, the metrics $d_{H}$ and $d_{S}$ are equivalent. For more background on the Hausdorff and symmetric difference metrics, see the article [ShephardWebster-1965] by Shephard and Webster. For a general reference on convex geometry, we refer the reader to the book [SchneiderBook] of Schneider.
Remark 2.1.
It was conjectured in [BH-2024] that $d_{H}$ and $\delta_{j}$ induce the same topology on $\mathcal{K}_{j}^{d}$. More specifically, in Open Question 2.2 of [BH-2024] they asked the following
Open Question 2.2 (uniform bound).
Let $j\in\{2,\dotsc,d-1\}$. If $\{K_{i}\}$ is a sequence in $\mathcal{K}_{j}^{d}$ that converges to $K\in\mathcal{K}_{j}^{d}$ with respect to $\delta_{j}$, does it follow that $\{K_{i}\}$ is uniformly bounded, that is, does there exist $R>0$ such that $\bigcup_{i\in\mathbb{N}}K_{i}\subset RB_{d}$?
For the special case $j=d$, this question was answered affirmatively by Shephard and Webster in [ShephardWebster-1965, Lemma 11], and for $j=1$ see [Vitale-1985, Theorem 3] and [Florian-1989, Theorem 2]. Theorem 1.1 gives a negative answer to this question for $2\leq j\leq d-1$.
3. Proofs
3.1. Proof of Theorem 1.1
Let $E:=\operatorname{span}\{e_{1},\ldots,e_{j}\}\subset\mathbb{R}^{d}$ and fix a $j$-dimensional convex body $K\subset E$ (so $K\in\mathcal{K}_{j}^{d}$). Choose $x_{0}\in\operatorname{relint}(K)$ (relative to $E$) and a unit vector $u\in E$. Shrinking $\varepsilon$ if necessary, we may assume that $x_{0}+\varepsilon B_{u^{\perp}\cap E}^{j-1}\subset K$. For parameters $L>0$ and $\varepsilon>0$, define the following thin cylinder inside $E$:
(2) $N(L,\varepsilon):=\left\{x_{0}+tu+v:\,t\in[0,L],\,v\in\varepsilon B_{u^{\perp}\cap E}^{j-1}\right\}\subset E.$
Here $B_{u^{\perp}\cap E}^{j-1}:=\{x\in u^{\perp}\cap E:\,\|x\|\leq 1\}$. Now set
(3) $K(L,\varepsilon):=\operatorname{conv}\left(K\cup N(L,\varepsilon)\right)\subset E.$
Note that $\dim K(L,\varepsilon)=j$ and $K(L,\varepsilon)\in\mathcal{K}_{j}^{d}$ for all $L,\varepsilon>0$.
Fix $H\in\operatorname{Gr}(d,j)$. Let $u_{H}:=(P_{H}u)/\|P_{H}u\|$ when $P_{H}u\neq 0$, and decompose $H$ as the direct sum $H=\mathbb{R}u_{H}\oplus E_{H}$ where $E_{H}:=(P_{H}u)^{\perp}\cap H$. Let $\pi_{H}$ denote the orthogonal projection along $u_{H}$. Then for every $y\in E_{H}$, the fiber $(P_{H}(K(L,\varepsilon)))_{y}\setminus(P_{H}K)_{y}$ is either empty or is a single segment of length at most $L\|P_{H}u\|$, and this can only occur when $y\in\pi_{H}(P_{H}(\varepsilon B_{u^{\perp}\cap E}^{j-1}))$. Hence, with $\ell_{H}:=\|P_{H}u\|$ we get
$\operatorname{vol}_{j}\left([P_{H}(K(L,\varepsilon))]\triangle(P_{H}K)\right)\leq L\ell_{H}\operatorname{vol}_{j-1}\left(\pi_{H}(P_{H}(\varepsilon B_{u^{\perp}\cap E}^{j-1}))\right).$
Therefore, integrating over $H$ we obtain
$\delta_{j}(K(L,\varepsilon),K)$ $\leq\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\int_{\operatorname{Gr}(d,j)}L\ell_{H}\operatorname{vol}_{j-1}(P_{H}(\varepsilon B_{u^{\perp}\cap E}^{j-1}))\,d\nu_{j}(H).$
The orthogonal projection is 1-Lipschitz on $H$ and thus does not increase the (j1)(j-1)-dimensional volume. Hence,
$\operatorname{vol}_{j-1}(P_{H}(\varepsilon B_{u^{\perp}\cap E}^{j-1}))\leq\operatorname{vol}_{j-1}(\varepsilon B_{u^{\perp}\cap E}^{j-1})=\operatorname{vol}_{j-1}(B_{u^{\perp}\cap E}^{j-1})\varepsilon^{j-1}=\operatorname{vol}_{j-1}(B_{j-1})\varepsilon^{j-1}.$
Thus, using also $\ell_{H}\leq 1$, we get
$\delta_{j}(K(L,\varepsilon),K)$ $\leq L\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\operatorname{vol}_{j-1}(B_{j-1})\varepsilon^{j-1}=C(d,j)L\varepsilon^{j-1}$
where $C(d,j):=\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\operatorname{vol}_{j-1}(B_{j-1})$.
Now choose a sequence $\{L_{i}\}$ with $L_{i}\to\infty$, and set $\varepsilon_{i}:=L_{i}^{-2}$. For any fixed $j\geq 2$, we have
$\delta_{j}(K(L_{i},\varepsilon_{i}),K)\leq C(d,j)L_{i}\varepsilon_{i}^{j-1}=C(d,j)L_{i}^{1-2(j-1)}=C(d,j)L_{i}^{3-2j}\to 0\quad\text{as}\quad i\to\infty.$
and $\sup_{x\in K(L_{i},\varepsilon_{i})}\|x\|\to\infty$ as $i\to\infty$. Thus, we have proved the first assertion of the theorem with $K_{i}:=K(L_{i},L_{i}^{-2})\in\mathcal{K}_{j}^{d}$.
To prove the second assertion, consider the point $x_{i}:=x_{0}+L_{i}u\in N(L_{i},\varepsilon_{i})\subset K(L_{i},\varepsilon_{i})$. Let
$R_{K}:=\sup\{\|x\|:\,x\in K\}<\infty.$
Observe that by the reverse triangle inequality,
$\|x_{i}\|=\|x_{0}+L_{i}u\|\geq\|L_{i}u\|-\|x_{0}\|=L_{i}-\|x_{0}\|,$
and, for any $x\in K$,
$\|x_{i}-x\|\geq\|x_{i}\|-\|x\|\geq L_{i}-\|x_{0}\|-R_{K}.$
Therefore,
$d_{H}(K(L_{i},\varepsilon_{i}),K)\geq\operatorname{dist}(x_{i},K)\geq L_{i}-\|x_{0}\|-R_{K}\longrightarrow\infty\quad\text{as}\quad i\to\infty.$
3.2. Proof of Theorem 1.2
Fix a $j$-dimensional subspace $E\in\mathrm{Gr}(d,j)$ and a unit vector $u\in E$. Given $x\in E$, $L>0$ and $\varepsilon>0$, define the following thin cylinder (a $j$-dimensional convex body inside $E$):
$N(x;L,\varepsilon):=\operatorname{conv}\left(x+[-L,L]u,x+\varepsilon B^{j-1}_{u^{\perp}\cap E}\right)\subset E,$
where $B^{j-1}_{u^{\perp}\cap E}$ is the Euclidean unit ball in $u^{\perp}\cap E$. For any $K\in\mathcal{K}^{d}_{j}$, set
$K^{+}(x;L,\varepsilon):=\operatorname{conv}(K\cup N(x;L,\varepsilon)).$
There exists a constant $C(d,j)>0$ (depending only on $d,j$) such that for all $K$, all $x\in E$, and all $L,\varepsilon>0$,
(4) $\delta_{j}(K^{+}(x;L,\varepsilon),K)\leq C(d,j)L\varepsilon^{j-1}.$
Indeed, for each $H\in\mathrm{Gr}(d,j)$, the $j$-dimensional volume in $H$ contributed by the added cylinder is bounded above by
$(\text{length of }P_{H}\text{-image of the segment})$ $\times((j-1)\text{-volume of the transverse cross-section})$
$\leq 2L\operatorname{vol}_{j-1}(B_{j-1})\varepsilon^{j-1},$
and we integrate this bound over $H$ (the projection is $1$-Lipschitz on $E$). Multiplying both sides of this inequality by the fixed factor $\big[\!\begin{smallmatrix}d\\ j\end{smallmatrix}\!\big]$ and setting $C(d,j):=2\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\operatorname{vol}_{j-1}(B_{j-1})$, we obtain (4).
Next, we construct a $\delta_{j}$-Cauchy sequence whose projections escape to infinity. Fix any $K_{0}\in\mathcal{K}^{d}_{j}$ and let $E$ and $u$ be as above. Choose a sequence $\{L_{m}\}_{m\geq 0}$ with $L_{m}\to\infty$. Define $\varepsilon_{m}>0$ by
(5) $\varepsilon_{m}^{j-1}L_{m}=\frac{2^{-(m+1)}}{C(d,j)}.$
We inductively define
$K_{m+1}:=\operatorname{conv}(K_{m}\cup N(x_{m};L_{m},\varepsilon_{m})),$
where $x_{m}:=x_{0}+T_{m}u$ and $T_{m}\to\infty$ is arbitrary (e.g., $T_{m}=m$). Then for a.e. $H$ with $\|P_{H}u\|>0$, we have $\|P_{H}x_{m}\|>\|P_{H}u\|T_{m}-\|P_{H}x_{0}\|\to\infty$ as $m\to\infty$. Hence, for any prescribed $R_{m}\to\infty$, we eventually have $P_{H}x_{m}\in H\setminus B_{H}(0,R_{m})$. In particular, the projected cylinder $P_{H}N(x_{m};L_{m},\varepsilon_{m})$ lies outside $B_{H}(0,R_{m})$ for all sufficiently large $m$. By (4) and (5), we have
(6) $\delta_{j}(K_{m+1},K_{m})\leq C(d,j)L_{m}\varepsilon_{m}^{j-1}\ =2^{-(m+1)}.$
Hence $\sum_{m}\delta_{j}(K_{m+1},K_{m})<\infty$, and $\{K_{m}\}$ is $\delta_{j}$-Cauchy.
Suppose by way of contradiction that $\{K_{m}\}$ converges in the metric $\delta_{j}$ to some $K_{\infty}\in\mathcal{K}^{d}_{j}$:
$\delta_{j}(K_{m},K_{\infty})=\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\int_{\operatorname{Gr}(d,j)}\operatorname{vol}_{j}((P_{H}K_{m})\triangle(P_{H}K_{\infty}))\,d\nu_{j}(H)\longrightarrow 0\quad\text{as}\quad m\to\infty.$
Then with $f_{m}(H):=\operatorname{vol}_{j}((P_{H}K_{m})\triangle(P_{H}K_{\infty}))$, this implies $f_{m}\to 0$ in $L^{1}(H)$. By the definition of $\delta_{j}$, since $f_{m}\to 0$ in $L^{1}(H)$ there exists a subsequence (still denoted $\{f_{m}\}$) such that for $\nu_{j}$-almost every $H\in\mathrm{Gr}(d,j)$,
(7) $f_{m}(H)=\operatorname{vol}_{j}((P_{H}K_{m})\triangle(P_{H}K_{\infty}))\longrightarrow 0\quad\text{as}\quad m\to\infty.$
Next, we show that for a.e. $H$, each new cylinder contributes a fixed positive $j$-dimensional volume outside $B_{H}(0,R_{m})=\{y\in H:\,\|y\|\leq R_{m}\}$. We will need the following lemma, proved below.
Lemma 3.1.
Fix $E\in\mathrm{Gr}(d,j)$ and a unit vector $u\in E$. For $\nu_{j}$-almost every $H\in\mathrm{Gr}(d,j)$, the map $P_{H}|_{E}:E\to H$ has full rank $j$, and the linear map
$T_{H}:=\bigl(\pi_{H}\circ P_{H}\bigr)\big|_{u^{\perp}\cap E}:\ u^{\perp}\cap E\longrightarrow E_{H},$
where $E_{H}:=\{x\in H:\langle x,u_{H}\rangle=0\}$ and $u_{H}:=\frac{P_{H}u}{\|P_{H}u\|}$, is an isomorphism with (j1)(j-1)-Jacobian $J_{j-1}(T_{H})>0$.
Fix an $H$ that satisfies Lemma 3.1, and set
$\ell_{H}:=\|P_{H}u\|>0,\qquad b(H):=J_{j-1}(T_{H})\operatorname{vol}_{j-1}(B_{j-1})>0,\qquad c(H):=2\ell_{H}b(H)>0.$
Since $P_{H}$ is linear and $T_{H}$ is an isomorphism, the $H$-projection of the cylinder contains a rectangular block,
$P_{H}N(x_{m};L_{m},\varepsilon_{m})\ \supset\ \Big(P_{H}x_{m}+[-L_{m},L_{m}]\cdot P_{H}u\Big)\ \times\ \pi_{H}\!\Big(P_{H}\big(\varepsilon_{m}B^{\,j-1}_{u^{\perp}\cap E}\big)\Big),$
whose $j$-dimensional volume is at least
(8) $\operatorname{vol}_{j}\big(P_{H}N(x_{m};L_{m},\varepsilon_{m})\big)\geq 2L_{m}\|P_{H}u\|\varepsilon_{m}^{j-1}J_{j-1}(T_{H})\operatorname{vol}_{j-1}(B_{j-1})=c(H)\varepsilon_{m}^{j-1}L_{m}.$
By (5), the right-hand side equals $c(H)C(d,j)^{-1}2^{-(m+1)}$.
Since $P_{H}K_{m}\subset B_{H}(0,\rho_{m})$ for some $\rho_{m}<\infty$, choose $R_{m}>\rho_{m}$. Thus, for all large enough $m$, $P_{H}N(x_{m};L_{m},\varepsilon_{m})\subset H\setminus B_{H}(0,R_{m})$ and hence is disjoint from $P_{H}K_{m}$. Thus, since $P_{H}K_{m+1}=\operatorname{conv}(P_{H}K_{m}\cup P_{H}N(x_{m};L_{m},\varepsilon_{m}))$, and since we choose $R_{m}$ so that $P_{H}K_{m}\subset B_{H}(0,\rho_{m})\subset B_{H}(0,R_{m})$, for $\nu_{j}$-a.e. $H\in\operatorname{Gr}(d,j)$, there exists $m_{0}\in\mathbb{N}$ such that for all $m\geq m_{0}$,
$\big(P_{H}K_{m+1}\setminus P_{H}K_{m}\big)\cap\big(H\setminus B_{H}(0,R_{m})\big)\supset P_{H}N(x_{m};L_{m},\varepsilon_{m}).$
and
(9) $\begin{split}\operatorname{vol}_{j}\Big(\big(P_{H}K_{m+1}\setminus P_{H}K_{m}\big)\cap\big(H\setminus B_{H}(0,R_{m})\big)\Big)&\geq\operatorname{vol}_{j}\big(P_{H}N(x_{m};L_{m},\varepsilon_{m})\big)\\ &\geq c(H)C(d,j)^{-1}2^{-(m+1)}.\end{split}$
In the final step, we will obtain a contradiction using $L^{1}(H)$-convergence to a bounded projection. Fix the above $H$. The sets $P_{H}K_{\infty}$ are bounded in $H$, so there exists $R>0$ such that $P_{H}K_{\infty}\subset B_{H}(0,R)$. From (7), for the subsequence (which we still denote by $\{K_{m}\}$), we have
$\operatorname{vol}_{j}\big((P_{H}K_{m})\triangle(P_{H}K_{\infty})\big)\longrightarrow 0\quad\text{as}\quad m\to\infty.$
In particular, for this fixed $R$, $\operatorname{vol}_{j}((P_{H}K_{m})\setminus B_{H}(0,R))\to 0$ as $m\to\infty$. But $R_{m}\to\infty$, so eventually $R_{m}>R$, and then (9) yields a uniform positive lower bound
$\operatorname{vol}_{j}((P_{H}K_{m+1})\setminus B_{H}(0,R))\geq c(H)C(d,j)^{-1}2^{-(m+1)}>0$
for infinitely many $m$. This contradicts the fact that $\operatorname{vol}_{j}\big((P_{H}K_{m})\setminus B_{H}(0,R)\big)\to 0$ as $m\to\infty$. Therefore, no $K_{\infty}\in\mathcal{K}^{d}_{j}$ can be the $\delta_{j}$-limit of $\{K_{m}\}$, so $(\mathcal{K}^{d}_{j},\delta_{j})$ is incomplete.
Proof of Lemma 3.1.
We first show the statement on the full rank of $P_{H}|_{E}$. For $v\in E\setminus\{0\}$, the set $\mathcal{A}_{v}:=\{H\in\mathrm{Gr}(d,j):\,H\subset v^{\perp}\}$ identifies with $\mathrm{Gr}(d-1,j)$, a proper smooth submanifold of strictly lower dimension. Hence, $\nu_{j}(\mathcal{A}_{v})=0$. Let $\mathbb{S}(E):=\{v\in E:\|v\|=1\}$ denote the unit sphere in $E$ centered at the origin. Since $\mathbb{S}(E)$ is compact, it contains a countable dense subset $Q\subset\mathbb{S}(E)$. For $q\in Q$, define
$\mathcal{A}_{q}:=\left\{H\in\mathrm{Gr}(d,j):\,H\subset q^{\perp}\right\}.$
Each $\mathcal{A}_{q}$ is naturally identified with $\mathrm{Gr}(d-1,j)$, so $\dim\mathcal{A}_{q}=j(d-j-1)<j(d-j)=\dim\mathrm{Gr}(d,j)$, and therefore $\nu_{j}(\mathcal{A}_{q})=0$. If $H\in\mathrm{Gr}(d,j)$ is such that $H\subset v^{\perp}$ for some nonzero $v\in E$, then $v/\|v\|\in\mathbb{S}(E)\cap H^{\perp}$, a nonempty closed subset of $\mathbb{S}(E)$. Since $Q$ is dense in $\mathbb{S}(E)$, there exists $q\in Q\cap H^{\perp}$. Thus $H\subset q^{\perp}$, so $H\in\mathcal{A}_{q}$. This yields
$\{H\in\operatorname{Gr}(d,j):\,\exists v\in E\setminus\{0\}\text{ with }H\subset v^{\perp}\}\subset\bigcup_{q\in Q}\mathcal{A}_{q},$
which is a countable union of $\nu_{j}$-null sets and thus is a $\nu_{j}$-null set. Thus, for $\nu_{j}$-a.e. $H$, $E\cap H^{\perp}=\{0\}$ and $P_{H}|_{E}$ is bijective.
Next, we show that $T_{H}$ is (j1)(j-1)-isomorphism on $u^{\perp}\cap E$. Fix any $H$ in which $P_{H}|_{E}$ has full rank, so $P_{H}u\neq 0$ and $u_{H}$ is defined. By the first part, this set has $\nu_{j}$ measure 1. If $T_{H}w=0$ for some $w\in u^{\perp}\cap E$, then $P_{H}w\in\mathrm{span}\{u_{H}\}$, say $P_{H}w=\lambda u_{H}=\lambda P_{H}u/\|P_{H}u\|$ for some $\lambda\in\mathbb{R}$. Then $P_{H}(w-\beta u)=0$ where $\beta=\lambda/\|P_{H}u\|$, and since $w-\beta u\in E$ and $P_{H}|_{E}$ is injective, we get $w-\beta u=0$. Hence $0=\langle u,w\rangle=\beta\langle u,u\rangle=\beta$, so $w=0$. Therefore, $T_{H}$ is injective. Since $\dim(u^{\perp}\cap E)=\dim(E_{H})=j-1$, $T_{H}$ is an isomorphism, and its (j1)(j-1)-Jacobian is positive. ∎
3.3. Proof of Theorem 1.3
Fix a $j$-dimensional subspace $E\in\operatorname{Gr}(d,j)$ with $2\leq j\leq d-1$, a unit vector $u\in E$, and a nonempty compact convex set $K_{0}\subset E$ with $\dim K_{0}=j$. For $L>0$, $\varepsilon>0$ and a base point $x_{0}\in K_{0}$, define the thin cylinder
$N(x_{0};L,\varepsilon):=\operatorname{conv}\left(x_{0}+[-L,L]u,x_{0}+\varepsilon B^{j-1}_{u^{\perp}\cap E}\right),$
and set $K^{+}(x_{0};L,\varepsilon):=\operatorname{conv}(K_{0}\cup N(x_{0};L,\varepsilon))$. As in the proof of Theorem 1.2, for all $L,\varepsilon>0$,
(10) $\delta_{j}(K^{+}(x_{0};L,\varepsilon),K_{0})\leq C(d,j)L\varepsilon^{j-1}.$
where $C(d,j):=2\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\operatorname{vol}_{j-1}(B_{j-1})>0$.
We will construct a $\delta_{j}$-Cauchy sequence with uniformly positive distance to $\varnothing$. Let $a_{0}:=\delta_{j}(K_{0},\varnothing)$, which is positive since $K_{0}\neq\varnothing$. Choose any sequence $L_{m}\to\infty$, and define
(11) $\varepsilon_{m}:=\left(\frac{2^{-(m+1)}a_{0}}{4C(d,j)L_{m}}\right)^{\frac{1}{j-1}}.$
Define inductively $K_{m+1}:=\operatorname{conv}(K_{m}\cup N(x_{m};L_{m},\varepsilon_{m}))$, where the choice of the base point $x_{m}\in E$ will be specified later. Then by (10) and (11),
(12) $\delta_{j}(K_{m+1},K_{m})\leq C(d,j)L_{m}\varepsilon_{m}^{j-1}=\frac{a_{0}}{4}\cdot 2^{-(m+1)}.$
Hence, $\sum_{m}\delta_{j}(K_{m+1},K_{m})\leq a_{0}/4$, so $\{K_{m}\}$ is $\delta_{j}$-Cauchy. Moreover, by the triangle inequality,
$\left|\delta_{j}(K_{m+1},\varnothing)-\delta_{j}(K_{m},\varnothing)\right|\leq\delta_{j}(K_{m+1},K_{m})\leq\frac{a_{0}}{4}\cdot 2^{-(m+1)},$
so $\{\delta_{j}(K_{m},\varnothing)\}$ is Cauchy in $\mathbb{R}$ and thus converges to some $\alpha\in[0,\infty)$. We now show that $\alpha>0$. Note that for every $m$,
$\delta_{j}(K_{m},\varnothing)\geq\delta_{j}(K_{0},\varnothing)-\delta_{j}(K_{0},K_{m})\geq a_{0}-\sum_{i=0}^{m-1}\delta_{j}(K_{i+1},K_{i})\geq a_{0}-\frac{a_{0}}{4}=\frac{3}{4}a_{0}>0.$
Passing to the limit, we thus obtain
(13) $\alpha=\lim_{m\to\infty}\delta_{j}(K_{m},\varnothing)\geq\frac{3}{4}a_{0}>0.$
Therefore, $\{K_{m}\}$ does not converge to $\varnothing$ in $(\mathcal{K}_{j}^{d}\cup\{\varnothing\},\delta_{j})$.
Let $\mathscr{G}\subset\operatorname{Gr}(d,j)$ denote the set of “good” subspaces in Lemma 3.1. Then by Lemma 3.1, $\nu_{j}(\mathscr{G})=1$. For fixed $H\in\mathscr{G}$, define
$\ell_{H}:=\|P_{H}u\|>0,\quad b(H):=J_{j-1}(T_{H})\operatorname{vol}_{j-1}(B_{j-1})>0,\quad c(H):=2\ell_{H}b(H)>0.$
We now choose the base points so that projected mass appears arbitrarily far from the origin. Let $\{R_{m}\}_{m\geq 0}$ be any sequence with $R_{m}\to\infty$. Since $P_{H}|_{E}$ is onto, for each $m$ we can choose the base point $x_{m}\in E$ so that $P_{H}x_{m}$ lies sufficiently far in the $u_{H}-$direction to guarantee
(14) $P_{H}N(x_{m};L_{m},\varepsilon_{m})\subset H\setminus B_{H}(0,R_{m}).$
Since the projection is linear and $T_{H}$ is an isomorphism, the projected cylinder contains a rectangular block in $H$ whose $j$-volume is bounded below by
(15) $\operatorname{vol}_{j}\big(P_{H}N(x_{m};L_{m},\varepsilon_{m})\big)\geq 2L_{m}\|P_{H}u\|\varepsilon_{m}^{j-1}J_{j-1}(T_{H})\operatorname{vol}_{j-1}(B_{j-1})=c(H)\varepsilon_{m}^{j-1}L_{m}.$
By (11), the right-hand side equals $\frac{c(H)a_{0}}{4C(d,j)}\cdot 2^{-(m+1)}$. Moreover, since $P_{H}K_{m}$ is contained in some ball $B_{H}(0,\rho_{m})$ with $\rho_{m}<\infty$ (as by continuity, projections of compact sets are compact), choosing $R_{m}>\rho_{m}$ makes $P_{H}N(x_{m};L_{m},\varepsilon_{m})$ disjoint from $P_{H}K_{m}$, and from (14) and (15) we get the lower bound
(16) $\operatorname{vol}_{j}\Big(\big(P_{H}K_{m+1}\setminus P_{H}K_{m}\big)\cap\big(H\setminus B_{H}(0,R_{m})\big)\Big)\ \geq\frac{c(H)a_{0}}{4C(d,j)}\cdot 2^{-(m+1)}>0.$
Finally, we show that the sequence $\{K_{m}\}$ has no limit in $\mathcal{K}_{j}^{d}$. Suppose by way of contradiction that $K_{m}\to K_{\infty}$ in $\delta_{j}$ for some $K_{\infty}\in\mathcal{K}_{j}^{d}$. Following along the same lines as before, there exists a subsequence (again, not relabeled) such that for $\nu_{j}$-a.e. $H$,
(17) $\operatorname{vol}_{j}\big((P_{H}K_{m})\triangle(P_{H}K_{\infty})\big)\longrightarrow 0\quad\text{as}\quad m\to\infty.$
Fix the good $H\in\mathscr{G}$ chosen above, so that (17) holds for these $H$. Since $K_{\infty}$ is compact, $P_{H}K_{\infty}\subset B_{H}(0,R)$ for some $R>0$. Since $R_{m}\to\infty$, for all sufficiently large $m$ we have $R_{m}>R$, and hence
$\operatorname{vol}_{j}\big((P_{H}K_{m})\setminus B_{H}(0,R)\big)\longrightarrow 0.$
But (16) yields, for infinitely many $m$,
$\operatorname{vol}_{j}\big((P_{H}K_{m+1})\setminus B_{H}(0,R)\big)\geq\frac{c(H)a_{0}}{4C(d,j)}\cdot 2^{-(m+1)}>0,$
a contradiction. Therefore, $\{K_{m}\}$ has no limit in $\mathcal{K}_{j}^{d}$.
To conclude the proof, note that by (13), the sequence does not converge to $\varnothing$, and by the last step, it does not converge to any $K_{\infty}\in\mathcal{K}_{j}^{d}$ either. Hence, $(\mathcal{K}_{j}^{d}\cup\{\varnothing\},\delta_{j})$ is not complete. ∎
Department of Mathematics & Computer Science, Longwood University,
201 High Street, Farmville, Virginia 23901
E-mail address:
hoehnersd@longwood.edu